# Learning Interpretable Hierarchical Dynamical Systems Models from Time Series Data

### v1

@@ 3.2 Training Method
For training, we used generalized teacher forcing (GTF) (Hess et al., 2023), a method specifically designed to address exploding gradients, which are known to occur when training DSR models on chaotic time series (Mikhaeil et al., 2022). GTF stabilizes training by linearly interpolating between RNN-generated and data-inferred control states in an optimal way (see Appx. A.1). This method was selected for its state-of-the-art performance across a range of DSR tasks (Hess et al., 2023). As in Hess et al. (2023), we trained directly on the observations, assuming $\bm{x}_{t}=h_{\bm{\theta}_{\text{obs}}^{(j)}}(\bm{z}_{t})=\bm{z}_{t}$ to be the identity mapping. We emphasize that GTF is only used to train the model, while at test time the models always run autonomously (i.e., without any feedback from actual observations).
Batching
During training, we used a batching strategy that ensures each gradient update incorporates data from all subjects. For datasets with a manageable number of subjects, we sampled training sequences $\bm{x}_{1...T_{\text{seq}}}^{(j)}$ from each subject, ensuring that updates reflect the collective dynamics across the entire population. For larger datasets, we sampled a random subset of subjects per batch. We further provided multiple sequences per subject within each minibatch to gather more robust statistics for the subject-specific feature vectors. Losses were then added for each subject-specific minibatch and combined to update both group-level parameters, $\bm{\theta}_{\text{group}}$, and subject-specific feature vectors $\bm{l}^{(j)}$. We found that a batch size between 416×4-16$\times$4 - 16 × the number of subjects led to the most robust training and performance (see Sect. A.1.2 and Figs. 8 & 9).
Differential learning rates
For training we used the RAdam (Liu et al., 2020) optimizer. To ensure successful training, we found it crucial to assign a significantly higher learning rate to the subject-specific feature vectors ($10^{-3}$) than to the group-level matrices ($10^{-4}$). This helped prevent numerical instabilities in the group-level matrices, which occurred with higher learning rates. It also prioritizes the incorporation of subject-specific information through the feature vectors, see Fig. 13.
Initialization
We used Xavier uniform initialization (Glorot and Bengio, 2010) for the group-level matrices, which is designed to keep the variance of the outputs $n_{out}$ of a layer roughly equal to the variance of its inputs $n_{in}$ by drawing weights from a uniform distribution in the interval $[-a,a]$, where $a=\sqrt{\frac{6}{n_{in}+n_{out}}}$. We reduced the weights further by a factor of $0.1$ to stabilize training. Additionally, we applied $L_{2}$ regularization to the group-level matrices to further enhance stability.
Learnable scaling matrices
Lastly, we introduced a learnable, subject-specific diagonal scaling matrix $\bm{\Sigma}$ in order to deal with data that is on different scales across subjects (e.g. for observations of the Lorenz-63 system in the cyclic vs. chaotic regime). This was incorporated by the use of a Gaussian negative log likelihood loss
$\mathcal{L}(\bm{x}_{t},\hat{\bm{x}}_{t})=-\frac{M}{2}\log|\bm{\Sigma}|-\frac{1}{2}(\bm{x}_{t}-\hat{\bm{x}}_{t})^{T}\bm{\Sigma}^{-1}(\bm{x}_{t}-\hat{\bm{x}}_{t})+const.$ (5)
in place of the MSE loss suggested in Hess et al. (2023) for training with GTF.
@@ 4 Results
In the following sections, we highlight various use cases of our hierarchical DSR framework.
@@ 4.1 Transfer Learning
To illustrate transfer learning, we used two popular DS benchmarks, where we sampled multivariate time series $\bm{X}^{(j)}$ with varying ground-truth parameters of the underlying ODE systems.
  • 1)
    Lorenz-63 data consisting of 64 ‘subjects’, simulated by 64 different parameter combinations for the Lorenz-63 model (Lorenz, 1963) of atmospheric convection (Eq. 13), with $\rho\in\{21,51,81,111\}$, $\sigma\in\{8,9,10,11\}$ and $\beta\in\{1,2,3,4\}$.
  • 2)
    Lorenz-96 data, based on a spatially extended version of the Lorenz model (Lorenz, 1996) with $N=10$ dimensions (Eq. 15), consisting of 20 subjects with $F\in\{10\cdot(n-1)+1\}_{n=1}^{20}$.
@@
We compared the performance of our framework to three other recent methods (see Appx. A.4 for details): First, as a baseline we tested an ensemble of individual shPLRNNs, using an otherwise identical training algorithm (a kind of ablation experiment, removing specifically the ‘hierarchical component’). Second, we employed LEarning Across Dynamical Systems (LEADS, Yin et al. (2021)), a framework that trains Neural ODEs for generalizing across DS environments by learning a shared dynamics model jointly with environment-specific models. Third, we trained context-informed dynamics adaptation (CoDA, Kirchmeyer et al. (2022)), an extension of LEADS where parameters of the combined and environment-specific models are drawn from a hypernetwork. As evidenced in Table 1, our hierarchical approach (hier-shPLRNN) considerably outperforms all other setups. In fact, competing methods were often not even able to correctly reproduce the long-term attractor dynamics (Appx. Fig. 12), while our approach successfully recovered different attractor topologies (Figs. 1 & 15).
@@ 4.2 Interpretability
We assessed the ability of the hierarchical inference framework to discover interpretable structure from the Lorenz-63 (Eq. 13) and Rössler system (Eq. 14). To this end, we again sampled relatively short time series of length $T=1000$ for 10 different values $\rho^{(j)}\in\{28\dots 80\}$ for the Lorenz-63, and $c^{(j)}\in\{3.8\dots 4.8\}$ for the Rössler, a range where its dynamics undergoes a bifurcation from limit cycle to chaotic dynamics. To reflect these 1-parameter-variation ground truth settings, we also chose $N_{\text{feat}}=1$ for the length of the subject-specific parameter vectors. After training, we found that the extracted feature values $\bm{l}^{(j)}$ were highly predictive of the ground truth values for $\rho^{(j)}$ and $c^{(j)}$, with a clearly linear relation (Fig. 2a, see Fig. 17 for the Lorenz-96). The observation that the model automatically inferred such a linear relationship is particularly noteworthy given that the DSR model’s piecewise-linear form (Eq. 4) profoundly differs from the polynomial equations defining the ground truth systems (Eqs. 13 and 14). Surprisingly, even when hier-shPLRNNs were trained with many more features ($N_{\text{feat}}=10$) than theoretically required, a principal component analysis (PCA) on the feature vectors revealed that a dominant part of the variation was captured by the first PC ($>85\%$ of variance), with negligible contributions beyond the third PC (Fig. 2b, blue curve). In contrast, attempts to extract low-dimensional structure from the parameters of individually trained models were unsuccessful, with variation distributed across many PCs (Fig. 2b, gray curve).
@@
We next examined a scenario where all three of the ground-truth parameters $\sigma,\rho$ and $\beta$ of the Lorenz-63 (Eq. 13) were jointly varied across a grid of $4\times 4\times 4=64$ subjects, as in Sect. 4.1, covering both periodic and chaotic regimes. While naturally, in this case, a scalar subject-specific feature did not capture the full variation in the dataset, we found that good reconstructions could already be achieved using $N_{\text{feat}}=3$, while using $N_{\text{feat}}=6$ features led to optimal performance (potentially indicating that for more complex, higher-dimensional scenarios some ‘disentangling’ may be required in feature space to retain linearity in $G_{\bm{\theta}_{\text{group}}}$, Eq. 2, despite the highly nonlinear relation of parameters to dynamics in the ground truth model). The extracted features remained highly interpretable, as illustrated in Fig. 3: $\rho$ strongly aligned with the first and $\beta$ with the second PC of the subject feature space, while variation in $\sigma$ appeared nested within steps of $\beta$ across PC2. The eight subjects on the left which do not fall into the $4\times 4$ grid with the others, correspond to parameter combinations that put the system into a non-chaotic, cyclic regime. These subjects do not only form a distinct group in the first two PCs, but also have significant non-zero third and fourth PCs (see Fig. 16). This observation helps explain why the algorithm benefits from additional features ($N_{\text{feat}}>3$), as these allow to represent different dynamical regimes as distinct regions in feature space, thereby further supporting its interpretability.
@@
After training, we found that the hier-shPLRNN became an almost perfect emulator of the data, with freely generated trajectories [^1] By freely generated we mean that these are not merely forward predictions, but new trajectories drawn from the generative DSR model using only a data-inferred initial condition. closely matching the ground truth pulse waves (Fig. 4a and Appx. Fig. 18 for further examples). This is particularly impressive given that the number of parameters of the hier-shPLRNN constituted less than $1\%$ of the total amount of training data points, and a comparatively low-dimensional (in relation to the number of biophysical parameters) feature vector ($N_{\text{feat}}=12$) was sufficient to capture individual differences between subjects. The model’s ability to achieve these reconstructions with so few parameters indicates that it extracted meaningful structure from the data, rather than merely memorizing it. Accordingly, the extracted feature vectors could predict the haemodynamic parameters via linear regression with high – and in all cases statistically significant (F-tests, $p<0.05$) – accuracy (mean $R^{2}=0.83$, Fig. 4b). [^2] Further, albeit small improvements in accuracy could be achieved by nonlinear regression via random forests (mean $R^{2}=0.87$). Two parameters with a particularly strong relationship ($R^{2}>0.97$) are in Fig. 4c. Moreover, the explained variances of the first five principal components of the feature space agreed well with that across the haemodynamic (biophysical) parameter space (Fig. 4d), indicating that the learned feature space accurately captured the pattern of variation in biophysical dynamics. Again, these results are highly non-trivial as the model was trained only on the arterial pulse waves without any knowledge of the biophysical ground truth parameters.
@@ 4.3 Few-Shot-Learning
To test ‘few-shot learning’, by which we mean here the ability to infer new models from only small amounts of data, we first trained a hier-shPLRNN on $S=9$ subjects simulated using the Lorenz-63 with $\rho^{(j)}_{\text{train}}=\{28,33.8,39.6,45.3,56.9,62.7,68.4,74.2,80\}$. We then generated new, short sequences $\bm{x}_{1...T_{\text{max}}}^{\text{test}}$ with $T_{\text{max}}=100$, using values of $\rho^{(j)}_{\text{test}}$ randomly sampled from the same interval $[28,80]$ that also contained the training data, and fine-tuned only a new feature vector $\bm{l}^{(j)}_{\text{test}}$ on this test set. The model closely approximated (interpolated) the true $\rho^{(j)}_{\text{test}}$ values from these single short sequences (Fig. 5a), and was even able to extrapolate to new $\rho^{(j)}_{\text{test}}$ outside the training range. Fig. 5b exemplifies how estimates become increasingly robust for longer sequences. This result is noteworthy as training an individual shPLRNN for the Lorenz-63 is challenging even with $1000$ time steps (see Table 1 &\&& Fig. 10). Even for sequences as short as $T_{\text{max}}=100$, the loss curves were often uni-modal and smooth around the true parameter values (Fig. 5c), as observed previously for training with sparse or generalized teacher forcing (Hess et al., 2023; Brenner et al., 2024b). This allowed gradient descent to converge rapidly, within $6$ seconds on a single $11^{\text{th}}$ Gen $2.30$ GHz Intel Core i7-11800H. It also enabled efficient use of $2^{\text{nd}}$-order, Hessian-based methods (Newton-Raphson), which converged within just $1$ second. Hence, even with minimal data fast and reliable estimation of previously unseen dynamics is feasible, extending into regimes beyond the training domain.
Figure 5: a: Predictions (by linear regression as above) for 25 $\rho^{(j)}_{\text{test}}$ values randomly drawn from the training range (green vertical lines) and several $\rho^{(j)}_{\text{test}}$ values extrapolated beyond the training range, on which only the scalar subject feature ($N_{\text{feat}}=1$) was fine-tuned on single short test sequences ($T_{\text{max}}=100$) (see Appx. A.1.5 on how estimates of standard deviation were obtained). b: Predictions of a $\rho^{\text{test}}$ as a function of training sequence length. c: GTF loss for a short training sequence ($T_{\text{max}}=100$) across different values of the scalar subject feature mapped onto $\rho$ via linear regression. Note the uni-modal loss centered on the true $\rho^{\text{test}}$ even for this short test set.
@@ 4.4 Experiments with human empirical data
In the previous sections, we directly linked extracted feature vectors to known ground truth parameters. However, in most real-world scenarios, ground truth parameters are unavailable and we only have indirect information like class labels, making feature interpretation more challenging. To assess the performance of our hierarchical framework in such settings, we trained it on human EEG data from subjects who were either healthy or experiencing epileptic seizures (Zhang et al., 2022; Andrzejak et al., 2001), as before in a completely unsupervised fashion (i.e., ignoring known class labels). Using Gaussian Mixture Models (GMMs) for clustering in the extracted DS feature space, we were able to recover the true class labels with an accuracy of $92.6\pm 1.2\%$ across ten runs, substantially outperforming several leading approaches for unsupervised feature extraction from time series, such as tsfresh, (Christ et al., 2018), Catch-22, (Lubba et al., 2019), ROCKET (Dempster et al., 2020), MiniRocket (Dempster et al., 2021), and attention-based convolutional autoencoders (Wang et al., 2023) (see Table 5). Fig. 6a shows that the two classes form distinct clusters clearly separated along the first PC of the training data, similar to the clustering of different dynamical regimes in Fig. 3, but unlike the other unsupervised techniques where classes were much less distinct.
Figure 6: a: First two principal components of the embedding spaces of the hier-shPLRNN and other unsupervised time series feature extraction approaches from Table 5. Ground truth labels are indicated by color (blue: epileptic, red: healthy) and predicted labels are indicated by markers (circle: predicted epileptic, cross: predicted healthy). The DS feature space of the hier-shPLRNN exhibits the clearest separation of epileptic and healthy subjects (as confirmed quantitatively). b: Left: Ground truth data for two correctly and one incorrectly classified subjects, colored by ground truth labels. Right: Model-generated trajectories, colored by predicted labels.
Fig. 6b further illustrates why mis-classifications happened: The subject shown in the bottom panel is dynamically much more similar to the one shown in the center than the one in the top panel, resulting in similar feature vectors for the center and bottom cases and hence assignment to the same class. An interesting question is why apparently the epileptic subject exhibits seemingly healthy activity, whether for instance there was a labeling error by the experimenter, or whether there is some artifact or electrode misplacement. Unfortunately, we do not have this information, but this case illustrates that even in situations where the class labels are (partially) known, our unsupervised methodology based on similarity in system dynamics may reveal interesting information. Taken together, these results demonstrate that our hierarchical framework can automatically discover dynamically relevant features, which allow for the identification of clinically relevant subgroups in an unsupervised way. In particular, our findings confirm that clustering according to DS properties may profoundly improve on ‘classical’ time series feature extraction, as conjectured.
@@
Finally, we tested our hierarchical DSR approach on a set of short fMRI time series from $2\times 10$ brain regions from 26 human subjects participating in a cognitive neuroscience experiment (Koppe et al., 2014; 2019), of which 10 with minimal irregularities or artifacts were selected. Based on PCA (Fig. 20c), we determined that a feature vector dimension of $N_{\text{feat}}=10$ was necessary to adequately capture the unique characteristics of the individual time series (Fig. 19a). For a left-out subject, reconstruction of the overall dynamics was feasible even when only fine-tuning the low-dimensional feature vector on unseen data (Fig. 19b). Since for this dataset we did not have any ground truth labels or classes, we sought to assess the robustness of feature extraction by computing the cosine similarity matrix across twenty training runs. Strong correlations ($r\approx 0.85\pm 0.05$) indicated that feature vectors were indeed robustly inferred. Clustering the features via GMMs revealed distinct groups that aligned with visual differences observed in the fMRI time series (Fig. 20).

### v2

@@ 3.2 Training Procedure
For training, we used generalized teacher forcing (GTF) (Hess et al., 2023), a SOTA DSR training method specifically designed to address exploding gradients, which are known to occur when training DSR models on chaotic time series (Mikhaeil et al., 2022). GTF stabilizes training by linearly interpolating between RNN-generated and data-inferred control states in an optimal way (see Appx. A.1). We emphasize that GTF is only used to train the model, while at test time the models always run autonomously. RAdam (Liu et al., 2020) was employed as the optimizer, with significantly higher learning rate for the subject-specific feature vectors ($10^{-3}$) than for the group-level matrices ($10^{-4}$), which was crucial to prevent numerical instabilities and prioritize incorporation of subject-specific information, see Fig. 19. Across subjects, a batching strategy was employed that ensured each gradient update incorporates data from all subjects, see Appx. A.1.4.
Note that our model, although multi-level, requires only fairly few hyper-parameters: the latent size $M$ (simply taken to be equal to the number of observations in all our tests), hidden size $L$ (for which $L=20\times M$ is a good default), GTF parameter $\alpha$ (which could be determined automatically, see Hess et al. (2023), but here was set to defaults), and the number of features $N_{\text{feat}}$; see Appx. A.1.11.
@@ 4.1 Transfer Learning
To illustrate transfer learning, we used two popular DS benchmarks, where we sampled multivariate time series $\bm{X}^{(j)}$ with varying ground-truth parameters of the underlying ODE systems: The Lorenz-63 model (Lorenz, 1963) of atmospheric convection (Eq. 15), simulated here with 64 different parameter combinations (‘subjects’) with $\rho\in\{21,51,81,111\}$, $\sigma\in\{8,9,10,11\}$ and $\beta\in\{1,2,3,4\}$; and the spatially extended Lorenz-96 model (Lorenz, 1996) with $N=10$ dimensions (Eq. 17), for which 20 subjects with $F\in\{10\cdot(j-1)+1\}_{j=1}^{20}$ were simulated.
@@
We compared the performance of our framework to three other recent methods (see Appx. A.4 for details): First, as a baseline we tested an ensemble of individual shPLRNNs, using an otherwise identical training algorithm (a kind of ablation experiment, removing specifically the ‘hierarchical component’). Second, we employed LEarning Across Dynamical Systems (LEADS, Yin et al. (2021)), a framework that trains Neural ODEs for generalizing across DS environments by learning a shared dynamics model jointly with environment-specific models. Third, we trained context-informed dynamics adaptation (CoDA, Kirchmeyer et al. (2022)), an extension of LEADS where parameters of the combined and environment-specific models are drawn from a hypernetwork. As evidenced in Table 1, our hierarchical approach (hier-shPLRNN) considerably outperforms all other setups. In fact, competing methods were often not even able to correctly reproduce the long-term attractor dynamics (Appx. Fig. 17), while our approach successfully recovered different attractor topologies (Figs. 2 & 21). Results stayed the same when vanilla RNNs trained by GTF or LSTMs were swapped in for the shPLRNN in our framework (Appx. A.1.6 & Tab. 3), though LSTMs performed quite poorly overall.
@@
Scaling and robustness
Figs. 12 & 13 show how performance scales with trajectory length $T_{\text{max}}$ and number of subjects $S$. While performance of the hier-shPLRNN is always substantially better than when models are trained individually, this is particularly evident for short time series, $T_{\text{max}}\leq 500$. For $T_{\text{max}}>500$ performance starts to plateau, but profound gains are still obtained by increasing $S$ up to $16$. Surprisingly, training times until a given performance level is reached actually decrease with the number of subjects $S$ (Fig. 14 & Appx. A.1.9). The reason presumably is that with higher $S$, the model has access to a larger portion of the underlying system’s parameter space, and can leverage this efficiently to speed up training on each single subject. Finally, robustness of discovered solutions across multiple training runs was very high, with feature vectors correlating by $r=0.84\pm 0.05$ on average across experiments, and as indicated by the $>10$-fold smaller MADs in Tab. 1 for the hier-shPLRNN compared to the other approaches ($\approx 5\%$ of the resp. medians), see also Fig. 11.
@@ 4.2 Interpretability
We assessed the ability of the hierarchical inference framework to discover interpretable structure from the Lorenz-63 (Eq. 15) and Rössler system (Eq. 16). To this end, we again sampled relatively short time series of length $T=1000$ for 10 different values $\rho^{(j)}\in\{28\dots 80\}$ for the Lorenz-63, and $c^{(j)}\in\{3.8\dots 4.8\}$ for the Rössler, a range where its dynamics undergoes a bifurcation from limit cycle to chaos. To reflect these 1-parameter-variation ground truth settings, we also chose $N_{\text{feat}}=1$ for the length of the subject-specific parameter vectors. After training, we found that the extracted feature values $\bm{l}^{(j)}$ were highly predictive of the ground truth control parameters $\rho^{(j)}$ and $c^{(j)}$, with a clearly linear relation (Fig. 3a, see Fig. 23 for the Lorenz-96). The observation that the model automatically inferred such a linear relationship is particularly noteworthy given that the DSR model’s piecewise-linear form (Eq. 4) profoundly differs from the polynomial equations defining the ground truth systems (Eqs. 15 and 16). Surprisingly, even when hier-shPLRNNs were trained with many more features ($N_{\text{feat}}=10$) than theoretically required, a principal component analysis (PCA) on the feature vectors revealed that a dominant part of the variation was captured by the first PC ($>85\%$ of variance), with negligible contributions beyond the third PC (Fig. 3b, blue curve). [^1] Interestingly, we observed that additional PCs reflected higher powers of the control parameters, i.e. PC2 was related to $\rho^{2}$ and PC3 to $\rho^{3}$ for the Lorenz-63. The same results were obtained with a very different DSR model, Markov Neural Operators incl. a dissipativity prior (Li et al., 2022), embedded into our approach (Fig. 18). In contrast, attempts to extract low-dimensional structure from the parameters of individually trained models were unsuccessful, with variation distributed across many PCs (Fig. 3b, gray curve).
@@
We next examined a scenario where all three of the ground-truth parameters $\sigma,\rho$ and $\beta$ of the Lorenz-63 (Eq. 15) were jointly varied across a grid of $4\times 4\times 4=64$ subjects, as in Sect. 4.1, covering both periodic and chaotic regimes. While naturally, in this case, a scalar subject-specific feature did not capture the full variation in the dataset, we found that good reconstructions could already be achieved using $N_{\text{feat}}=3$, while using $N_{\text{feat}}=6$ features led to optimal performance. The extracted features remained highly interpretable, as illustrated in Fig. 4: $\rho$ strongly aligned with the first and $\beta$ with the second PC of the subject feature space, while variation in $\sigma$ appeared nested within steps of $\beta$ across PC2. The eight subjects on the left which do not fall into the $4\times 4$ grid with the others, correspond to parameter combinations that put the system into a non-chaotic, cyclic regime. These subjects do not only form a distinct group in the first two PCs, but also have significant non-zero third and fourth PCs (see Fig. 22). This observation helps explain why the algorithm benefits from additional features ($N_{\text{feat}}>3$), as these allow to represent different dynamical regimes as distinct regions in feature space, thereby further supporting its interpretability.
Fig. 28 further demonstrates that this semantic structuring of the feature spaces is observed even when the hier-shPLRNN is trained simultaneously on different dynamical regimes from three completely different DS, the Lorenz, the Rössler and the Chua system. The different DS are widely separated in feature space, and are further segregated according to dynamical regime within a given DS.
@@
Capturing non-stationarity with time-dependent features
One interesting use case could be if the different systems are functionally related, e.g., reflect different temporal snapshots of a non-autonomously evolving DS with time-dependent control parameter(s). In Fig. 16 we simulated such a scenario for the Lorenz-63 with $\rho(t)\in[28,68]$ slowly drifting across time. Co-training an additional function $\tilde{\bm{l}}^{(j)}_{t}=\text{MLP}(\bm{l}^{(j)},t)$ resulted in an explicity time-dependent DSR model which almost perfectly captured the temporal drift in $\rho(t)$ (linear $R^{2}=0.995$), confirming that our approach could in principle easily be amended to handle such situations.
@@
After training, we found that the hier-shPLRNN became an almost perfect emulator of the data, with freely generated trajectories [^2] By freely generated we mean that these are not merely forward predictions, but new trajectories drawn from the generative DSR model using only a data-inferred initial condition. closely matching the ground truth pulse waves (Fig. 5a and Appx. Fig. 24 for further examples). This is particularly impressive given that the number of parameters of the hier-shPLRNN constituted less than $1\%$ of the total amount of training data points, and a comparatively low-dimensional (in relation to the number of biophysical parameters) feature vector ($N_{\text{feat}}=12$) was sufficient to capture individual differences between subjects. The model’s ability to achieve these reconstructions with so few parameters indicates that it extracted meaningful structure from the data, rather than merely memorizing it. Accordingly, the extracted feature vectors could predict the haemodynamic parameters via linear regression with high – and in all cases statistically significant (F-tests, $p<0.05$) – accuracy (mean $R^{2}=0.83$, Fig. 5b). [^3] Further, albeit small, improvements in accuracy could be achieved by nonlinear regression via random forests (mean $R^{2}=0.87$). Two parameters with a particularly strong relationship ($R^{2}>0.97$) are in Fig. 5c. Moreover, the explained variances of the first five principal components of the feature space agreed well with that across the haemodynamic (biophysical) parameter space (Fig. 5d), indicating that the learned feature space accurately captured the pattern of variation in biophysical dynamics. Again, these results are highly non-trivial as the model was trained only on the arterial pulse waves without any knowledge of the biophysical ground truth parameters.
@@ 4.3 Few-Shot-Learning
To test ‘few-shot learning’, by which we mean here the ability to infer new models from only small amounts of data, we first trained a hier-shPLRNN on $S=9$ subjects simulated using the Lorenz-63 with $\rho^{(j)}_{\text{train}}=\{28,33.8,39.6,45.3,56.9,62.7,68.4,74.2,80\}$. We then generated new, short sequences $\bm{x}_{1...T_{\text{max}}}^{\text{test}}$ with $T_{\text{max}}=100$, using values of $\rho^{(j)}_{\text{test}}$ randomly sampled from the same interval $[28,80]$ that also contained the training data, and fine-tuned only a scalar new feature $\bm{l}^{(j)}_{\text{test}}$ on this test set. The model closely approximated (interpolated) the true $\rho^{(j)}_{\text{test}}$ values from these single short sequences (Fig. 6a), and was even able to extrapolate to new $\rho^{(j)}_{\text{test}}$ outside the training range. Fig. 6b exemplifies how estimates become increasingly robust for longer sequences. This result is noteworthy as training an individual shPLRNN for the Lorenz-63 is challenging even with $1000$ time steps (see Table 1 &\&& Fig. 12). Even for sequences as short as $T_{\text{max}}=100$, the loss curves were often uni-modal and smooth around the true parameter values (Fig. 6c), as observed previously for training with sparse or generalized teacher forcing (Hess et al., 2023; Brenner et al., 2024). This allowed gradient descent to converge rapidly, within $6$ seconds on a single $11^{\text{th}}$ Gen $2.30$ GHz Intel Core i7-11800H. It also enabled efficient use of $2^{\text{nd}}$-order, Hessian-based methods, which converged within just $1$ second. Hence, even with minimal data fast and reliable estimation of previously unseen dynamics is feasible, extending into regimes beyond the training domain.
Figure 6: a: Predictions (by linear regression as above) for 25 $\rho^{(j)}_{\text{test}}$ values randomly drawn from the training range (green vertical lines) and several $\rho^{(j)}_{\text{test}}$ values extrapolated beyond the training range, on which only the scalar subject feature ($N_{\text{feat}}=1$) was fine-tuned on single short test sequences ($T_{\text{max}}=100$) (see Appx. A.1.10 on how estimates of standard deviation were obtained). b: Predictions of a $\rho^{\text{test}}$ as a function of training sequence length. c: GTF loss for a short training sequence ($T_{\text{max}}=100$) across different values of the scalar subject feature mapped onto $\rho$ via linear regression. Note the uni-modal loss centered on the true $\rho^{\text{test}}$ even for this short test set.
@@ 4.4 Experiments with human empirical data
In the previous sections, we directly linked extracted feature vectors to known ground truth parameters. However, in most real-world scenarios, ground truth parameters are unavailable and we only have indirect information like class labels, making feature interpretation more challenging. To assess the performance of our hierarchical framework in such settings, we trained it on human EEG data from subjects who were either healthy or experiencing epileptic seizures (Zhang et al., 2022; Andrzejak et al., 2001), as before in a completely unsupervised fashion (i.e., ignoring known class labels). Using Gaussian Mixture Models (GMMs) for clustering in the extracted DS feature space, we were able to recover the true class labels with an accuracy of $92.6\pm 1.2\%$ across ten runs, substantially outperforming several leading approaches for unsupervised feature extraction from time series, such as tsfresh, (Christ et al., 2018), Catch-22, (Lubba et al., 2019), ROCKET (Dempster et al., 2020), MiniRocket (Dempster et al., 2021), and attention-based convolutional autoencoders (Wang et al., 2023) (see Table 7). Fig. 7a shows that the two classes form distinct clusters clearly separated along the first PC of the training data, similar to the clustering of different dynamical regimes in Fig. 4, but unlike the other unsupervised techniques where classes were much less distinct. These findings confirm that clustering according to DS properties may profoundly improve on ‘classical’ time series feature extraction, as conjectured. Fig. 7b further illustrates why mis-classifications happened.
Figure 7: a: First two principal components of the embedding spaces of the hier-shPLRNN and other unsupervised time series feature extraction approaches from Table 7. Ground truth labels are indicated by color (blue: epileptic, red: healthy) and predicted labels are indicated by markers (circle: predicted epileptic, cross: predicted healthy). The DS feature space of the hier-shPLRNN exhibits the clearest separation of epileptic and healthy subjects (as confirmed quantitatively). b: Left: Ground truth data for two correctly and one incorrectly classified subjects, colored by ground truth labels. Right: Model-generated trajectories, colored by predicted labels. Note that the incorrectly classified subject is indeed dynamically much more similar to the ‘healthy’ than the ‘epileptic’ EEG.
@@
We further tested our hierarchical DSR approach on a set of short fMRI time series from $2$ (hemispheres) ×$\times 10$ distinct brain regions from 26 human subjects participating in a cognitive neuroscience experiment (Koppe et al. (2014), see Appx. A.3), of which 10 with minimal irregularities or artifacts were selected. Based on PCA (Fig. 26c), we determined that a feature vector dimension of $N_{\text{feat}}=10$ was necessary to adequately capture the unique characteristics of the individual time series (Fig. 25a). For a left-out subject, reconstruction of the overall dynamics was feasible even when only fine-tuning the low-dimensional feature vector on unseen data (Fig. 25b). Since for this dataset we did not have any ground truth labels or classes, we sought to assess the robustness of feature extraction by computing the cosine similarity matrix across twenty training runs. Strong correlations ($r\approx 0.85\pm 0.05$) indicated that feature vectors were indeed robustly inferred. Clustering the features via GMMs revealed distinct groups that aligned with visual differences observed in the fMRI time series (Fig. 26).
Finally, Fig. 27 illustrates that our hierarchical DSR approach also works for more ‘classical’ time series data (electricity consumption), beyond the domain of scientific applications.